Determine the concentration of diluted solution

Question:

Calculate the concentration of a solution prepared by diluting 0.058 L of 0.28 M SiF6 solution to a volume of 0.075 L.

Your answer should have two significant figures.
Answer #1

Number of moles = molarity * volume of solution in L

Number of moles = 0.28 * 0.058 = 0.016 mole

concentration of solution = number of moles / volume of solution in L

concentration of solution = 0.016 / 0.075 = 0.21 M

Therefore, concentration of solution = 0.21 M